Tyrell Co. entered into the following transactions involving short-term liabilities in 2016 and 2017 2016 Apr. 20 Purchased $37,000 of merchandise on credit from Locust, terms n/30. Tyrell uses the perpetual May 19 Replaced the April 20 account payable to Locust with a 90-day, S35.000 note bearing 8% annual July 8 Borrowed $69,000 cash from NBR Bank by signing a 120-day, 10% interest-bearing note with a face inventory system. interest along with paying $2,000 in cash. value of $69, 000. ?.. Paid the amount due on the note to Locust at the maturity date. Paid the amount due on the note to NBR Bank at the maturity date. Nov. 28 Borrowed $21,000 cash from Fargo Bank by signing a 60-day, 7% interest-bearing note with a face Dec. 31 Recorded an adjusting entry for accrued interest on the note to Fargo Bank 2017 value of $21,000. 7 Paid the amount due on the note to Fargo Bank at the maturity date. Required 1. Determine the maturity date for each of the three notes described NBR Bank Fargo Bank Matuity date
Ch (The following information applies to the questions displayed below Tyrell Co. entered into the following transactions involving short-term liabilities in 2016 and 20 2016 Apr. 20 Purchased $37,000 of merchandise May 19 Replaced the April 20 account payable to Locust with a 90-day, $35,000 note bearing 8% annual July 8 Borrowed $69,000 cash from NBR Bank by signing a 120-day, 10% interest-bearing note with a face on credit from Locust, terms n/30. Tyrell uses the perpetual inventory system. interest along with paying $2,000 in cash. value of $69,000 2 Paid the amount due on the note to Locust at the maturity date. Nov. 28 Borrowed $21,000 cash from Fargo Bank by signing a 60-day, 7% interest-bearing note with a face Dec. 31 Recorded an adjusting entry for accrued interest on the note to Fargo Bank. 2017 Paid the amount due on the note to NBR Bank at the maturity date. value of $21,000. Paid the amount due on the note to Fargo Bank at the maturity date. 2. Determine the interest due at maturity for each of the three notes. (Do not round your intermediate calculations. Use 360 days a Principal x Ratex Time Interest < Prey 11 12c13 14 of 14 Next >
Tyrell Co. entered into the following transactions involving short-term liabilities in 2016 and 2017 2016 Apr. 20 Purchased $37, 000 of merchandise on credit from Locust, terms n/30. Tyrell uses the perpetual May 19 Replaced the April 20 account payable to Locust with a 90-day, $35,000 note bearing 8% annual July 8 Borrowed $6,000 cash from NBR Bank by signing a 120-day, 10% interest-bearing note with a face inventory system. interest along with paying $2,000 in cash. value of $69, 000 一?-Paid the amount due on the note to Locust at the maturity date. Nov. 28 Borrowed $21,000 cash from Fargo Bank by signing a 60-day, 7% interest-bearing note with a face Dec. 31 Recorded an adjusting entry for accrued interest on the note to Fargo Bank. 2017 Paid the amount due on the note to NBR Bank at the maturity date. value of $21, 000 ? Paid the amount due on the note to Fargo Bank at the maturity date. 3. Determine the interest expense to be recorded in the adjusting entry at the end of 2016. (Do not round your intermediate calculations. Use 360 days a year.) Year end accrual required for Fargo Banlk Principal Ratex TimeInterest
Answer #1
1) Maturity date
locust NBR fargo
date of the note 19-May 8-Jul 28-Nov
term of note 90 120 60
maturity date 17-Aug 5-Nov 27-Jan
2) interest due at maturity
principal * Rate * time = interest
locust 35,000 * 8% * 90/360 = 700
NBR 69,000 * 10% * 120/360 = 2300
Fargo 21,000 * 7% * 60/360 = 245
3) Amount in adjusting entry
42,000*8%*33/360
135
Fargo Bank
principal * Rate * time = interest
interest to be acccrued in 2016 21,000 * 7% * 33/360 = 135
4) interest expense to be recorded in 2017
110
principal * Rate * time = interest
interest to recorded in 2018 21,000 * 7% * 27/360 = 110
Journal entries
Date Accounting titles & Explanations Debit Credit
2016
20-Apr inventory 37,000
Accounts payable 37,000
19-May Accounts payable 37,000
cash 2,000
notes payable 35,000
8-Jul Cash 69,000
notes payable 69,000
17-Aug notes payable 35,000
interest expense 700
cash 35,700
5-Nov notes payable 69,000
interest expense 2,300
cash 71,300
28-Nov Cash 21,000
notes payable 21,000
31-Dec interest expense 135
interest payable 135
2017
27-Jan notes payable 21,000
interest payable 135
interest expense 110
cash 21,245